how do you graph the linear function 8x + 5y = 600
Answer by TimothyLamb(4379)

You need two points on the line in order to graph it.
First get the equation in slope-intercept form: y = mx + b
---
8x + 5y = 600
5y = -8x + 600
y = (-8/5)x + (600/5)
---
Plug-in two different x values in the equation to get their y values, like so:
Point p1, use x=0:
y = (-8/5)(0) + (600/5)
y = (600/5)
y = 120
p1 = (0, 120)
---
Point p2, use x=50 (using x=50 makes y a whole number):
y = (-8/5)(50) + (600/5)
y = (-400/5) + (600/5)
y = (200/5)
y = 40
p2 = (50, 40)
---
Now draw a line between points p1 and p2 to graph the equation.
The line extends infinitely beyond both points.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php